Which of the following statements are true, and which are false? Give reasons for your answers. i) For any ring R and . a, b R, (a b) a 2ab b 2 2 2 ∈ + = + + ii) Every ring has at least two elements. iii) If R is a ring with identity and I is an ideal of R, then the identity of R I is the same as the identity of R. iv) If f : R → S is a ring homomorphism, then it is a group homomorphism from (R, +) to (S, +). v) If R is a ring, then any ring homomorphism from R ×R into R is surjective


1
Expert's answer
2021-05-07T09:53:38-0400

(i) True: Suppose R is a commutative ring. Then "ab=ba" for every "a, b ∈ R, \text{ so } −ab + ba = 0,"

so by the first equation above, "(a + b)(a \u2212 b) = a^2 + 0 \u2212 b^2 = a^2 \u2212 b^2" .


(ii) False: There is a unique ring with one element called the zero ring. It is a really bad idea to adopt the convention that the zero ring isn't a ring; without it the category of rings fails to have a terminal object and the tensor product of two rings (which is the coproduct for commutative rings) fails to exist in general.


(iii) True:As with subgroups, we call ideal I proper in R if I = 0 and I= R. Noticethat if R has identity (or “unity”) R and if R is in ideal I then I = R. In fact,if u is a unit in R and u ∈ I, then R = u−1 ∈ I and so I = R. If R ∈ thenI = R. So we can say that a left (or right) nonzero ideal I of ring R with unity Ris proper if and only if I contains no units of R. So a division ring (in which everynonzero element is a unit) has no proper left (or right) ideals.


(iv) False:Let ϕ_(R) = s. Observe that "s = \u03d5(R) = \u03d5(R.R) = \u03d5(R) \u00b7 \u03d5(R) = s^2."

Then we have "s = s^2" , and equivalently "s(1 \u2212 s) = 0" . Note that s − 1 = 0 by construction.

Now suppose that "\u03d5(R) = 0" , then, for any r in R "\u03d5(r) = \u03d5(r \u00b7 R) = \u03d5(r)\u03d5(R) = \u03d5(r) \u00b7 0 = 0"

and thus, "\u03d5" is the zero homomorphism, a contradiction. We conclude that "s(s \u2212 1) = 0"

holds for s = 0. Hence R is not surjective.